Three masses are located in the corners of an equillateral triangle. Find the magnitude and direction graitational field at the center of the triangle. Given:m1=22, m2=340,m3=340


1
Expert's answer
2021-04-26T17:11:51-0400

Answer

Assume that side of equilateral triangle

Is L.

Gravitational field at center of equilateral triangle is given by

"\vec{g}=G\frac{m_1}{(\frac{l}{\sqrt{3}})^2}\hat{j}-G\frac{m_2}{(\frac{l}{\sqrt{3}})^2}cos60°\hat{j}+G\frac{m_2}{(\frac{l}{\sqrt{3}})^2}sin60°\hat{i}-G\frac{m_3}{(\frac{l}{\sqrt{3}})^2}cos60°\hat{j}-G\frac{m_3}{(\frac{l}{\sqrt{3}})^2}sin60°\hat{j}"